# Practical hybrid decoding scheme for parity-encoded spin systems

### v3

@@ Abstract
We propose a practical hybrid decoding scheme for the parity-encoding architecture. This architecture was first introduced by N. Sourlas as a computational technique for tackling hard optimization problems, especially those modeled by spin systems such as the Ising model and spin glasses, and reinvented by W. Lechner, P. Hauke, and P. Zoller to develop quantum annealing devices. We study the specific model, called the SLHZ model, aiming to achieve a near-term quantum annealing device implemented solely through geometrically local spin interactions. Taking account of the close connection between the SLHZ model and a classical low-density-parity-check code, two approaches can be chosen for the decoding: (1) finding the ground state of a spin Hamiltonian derived from the SLHZ model, which can be achieved via stochastic decoders such as a quantum annealer or a classical Monte Carlo sampler; (2) using deterministic decoding techniques for the classical LDPC code, such as belief propagation and bit-flip decoder. The proposed hybrid approach combines the two approaches by applying bit-flip decoding to the readout of the stochastic decoder based on the SLHZ model. We present simulations demonstrating that this approach can reveal the latent potential of the SLHZ model, realizing soft-annealing concept proposed by Sourlas.
@@ I Introduction
The spin glass model is a fundamental concept in physics and mathematics, primarily used to study magnetism and phase transitions in statistical mechanics. A close connection between spin glass models and combinatorial optimization problems (COPs) is widely recognized and used as a tool for solving COPs [1]. The key idea is that the energy of the random Ising Hamiltonian can be viewed as a cost function whose ground state corresponds to a solution of the COP. Probabilistic simulation of Ising spin dynamics, such as simulated annealing (SA) [2] and quantum annealing (QA) [3], is often used to find the ground state. The spin glass model has applications in various industrial fields, including routing, scheduling, planning, decision-making, transportation, and telecommunications. A similar connection has been established between the spin glass model and the classical error-correcting codes (ECCs). N. Sourlas formulated the decoding of the classical ECCs in terms of Bayesian inference [4]. He showed that the decoding can be expressed as a search for the ground state of the random Ising Hamiltonian. His study implies that decoding classical ECCs can be viewed as a COP, with the associated random Ising Hamiltonian corresponding to the cost function to be minimized (maximized).
@@
The complexity class of the COPs is known to be NP-hard. When we use the spin glass models to tackle the COPs, the presence of many local minima separated by large energy barriers is a primary problem. The simulation algorithm gets trapped in those minima. To tackle this problem, Sourlas noted that two mathematically equivalent Hamiltonians can be derived for an ECC: a Hamiltonian given in terms of the original information source and an extended Hamiltonian given in terms of the information source and additional parity information [5, 6, 7]. The extended Hamiltonian involves extra parity-constraint terms. He proposed soft-annealing concept, which uses the extended Hamiltonian as a new approach to solve hard COPs [8]. The key idea is that one can circumvent barriers and accelerate the algorithm’s dynamics by enlarging the space on which the problem is defined.
@@
On the other hand, the QA is expected to resolve the above problem through adiabatic quantum computing [9]. However, to apply a QA device to solve the COP universally, we need to simulate a fully connected graph model in a scalable manner using Ising spin hardware. The soft-annealing approach is also helpful for this purpose. W. Lechner, P. Hauke, and P. Zoller (LHZ) proposed utilizing this approach to embed a random Ising model with all-to-all connectivity within an enlarged Ising system with geometrically local interactions only [10]. They proposed the parity-encoding (PE) architecture and its concrete realization, which we refer to as the SLHZ model in this paper. In this model, Ising spins are arranged on a plane. They are connected by four-body interactions in a lattice-like manner, which serve as a penalty that constrains the parities of spins. Both the local fields and couplings in the original spin glass model are mapped to local fields that act on each spin in the SLHZ model. This architecture is particularly attractive for realizing near-term quantum annealing (QA) devices based on superconducting quantum bits [11, 12, 13, 14] because it is highly compatible with on-chip electronic circuitry [15, 16, 17, 18, 19]. LHZ also claimed that the SLHZ model is a programmable, scalable, and robust physical platform for QA. Notably, Sourlas was the first researcher to propose soft annealing using the SLHZ model and recognize that this corresponds to the decoding of associated classical ECCs [8].
@@
Later, F. Pastawski and J. Preskill [20] studied the error correcting capability of the SLHZ model. They noted that the soft annealing usng the SLHZ model can be viewed as decoding of associated classical low-density parity-check (LDPC) code. Based on this observation, they applied the belief-propagation (BP) algorithm, a standard algorithm for LDPC codes [21], to decode the noisy readout of the SLHZ system [22, 23]. They found that one can obtain an error-free state from the noisy readout of the SLHZ model with a high probability if errors are generated by independent and identically distributed (i.i.d.) random spin-flip noise.
@@
To date, however, there has been little investigation focusing on whether the realistic classical decoding technique can fully exploit the potential of the SLHZ model. For example, T. Albash, W. Vincl, and D. Lidar discussed whether several decoding algorithms may boost the performance of the QA device [24]. They reported that simple majority-vote decoding (MVD) does not improve the performance of QA. Although they found that more sophisticated minimum-weight decoding (MWD) can boost performance, it is NP-hard and an unrealistic approach for post-readout decoding of QA devices (please refer to Ref. [24] and the Appendix for details). In the context of developing near-term QA, it is preferable to devise a more practical decoding strategy tailored for the SLHZ model that can exploit its potential.
@@
This paper demonstrates that realistic classical decoding techniques for LDPC codes can unveil potential in the SLHZ model that has not been recognized to date. We propose a practical and straightforward post-readout decoding algorithm tailored to the SLHZ model. The proposed algorithm is an iterative hard-decision decoding algorithm based on majority voting of generalized syndromes, known as Gallager’s bit-flipping (BF) algorithm, which was derived in the context of classical LDPC codes [22, 23]. Although the MWD is NP-hard, the BF algorithm is P, similar to the BP algorithms. Moreover, the BF algorithm requires less computational effort and processing time than the BP algorithm and can be implemented in digital logic. To demonstrate the BF algorithm, we performed two simulations. First, assuming i.i.d. spin-flip noise, we demonstrated that our BF decoding algorithm can recover the noisy output of the SLHZ model, eliminating spin-flip errors with performance comparable to that of the BP algorithm. Second, we performed classical simulations of stochastically sampled spin readouts from the SLHZ model using a Markov chain Monte Carlo (MCMC) sampler. This analysis allows us to investigate tolerance to leakage errors arising from dynamical and thermal excitations during the sampling. We show evidence that the BF decoding algorithm can efficiently eliminate spin-flip errors if we make the four-body penalty constraint sufficiently weak. This simulation suggests that a hybrid decoding approach, combining two algorithms — MCMC sampling followed by BF decoding — can be applied to a broader range of spin-flip noise and mitigate the computational overhead of SLHZ systems. Our results are consistent with earlier work by Albash et al., who employed simulated quantum annealing to simulate the readout of the QA device based on the SLHZ model (hereafter referred to as the SLHZ-based QA device) when BP decoding was used. We try to understand this result by comparing our algorithm with several extended BF decoding algorithms. According to our analysis, this is a consequence of the combination of the Hamiltonian of the SLHZ system and the properties of BF or BP decoding. We believe that our BF decoding is promising for realizing a near-term SLHZ-based QA device.
@@
This paper is organized as follows. Section II explains the PE architecture, the SLHZ model, and the close connection between classical LDPC codes and the SLHZ model. Section III describes classical decoding of the SLHZ model based on Bayesian inference. We propose the BF algorithm as a simple, straightforward decoding algorithm tailored for the SLHZ model. Section IV demonstrates the proposed BF algorithm under the i.i.d. and non-i.i.d. spin-flip noise, and shows the performance. In Sec.V, we compare several BF decoding algorithms and discuss why our two-stage hybrid decoding strategy tolerates errors originating from non-i.i.d. spin-flip noise. We explain why two-stage hybrid decoding outperforms the two algorithms individually. Section VI concludes this paper. We also discuss the relevance and consistency of our study to the earlier work by Albash et al. [24] in the Appendix A . Supplementary result is also provided in the Appendix B .
II Error correcting codes and parity-encoding architecture
@@ II.1 Notion
Suppose that a binary source-word $\tilde{M}$ is first encoded into a binary code-word $\tilde{C}$ using some ECC. The code-word is modulated into the physical signal $\tilde{S}$ and transmitted. The signal is affected by noise during transmission through a transmission channel. Let $\tilde{M}$ and $\tilde{C}$ be $K$ bits and $N_{v}\left(>K\right)$ bits, respectively, and denote them by the binary vectors $\boldsymbol{\bar{Z}}=(\bar{Z}_{1},\ldots,\bar{Z}_{K})\in\{0,1\}^{K}$ and $\bar{\boldsymbol{z}}=\left(\bar{z}_{1},\ldots,\bar{z}_{N_{v}}\right)\in\{0,1\}^{N_{v}}$, respectively. A linear code is defined by a one-to-one map from the set $\{\bar{\boldsymbol{Z}}\}$ of $2^{K}$ source-words $\tilde{M}$ of length $K$ to the set $\{\bar{\boldsymbol{z}}\}$ of $2^{K}$ code-words $\tilde{C}$ of length $N_{v}$. They are specified by a generating matrix $\boldsymbol{G}_{K\times N_{v}}$ or a generalized parity check matrix $\boldsymbol{H}_{N_{c}\times N_{v}}$ satisfying $\boldsymbol{G}\boldsymbol{H}^{T}=\boldsymbol{0}_{K\times N_{c}}\>\left(\mathrm{mod}\>2\right)$. Here $\boldsymbol{G}$ and $\boldsymbol{H}$ are binary matrices (i.e., their elements are 0 or $1$) and $\mathrm{mod}\>2$” denotes that the multiplication is modulo two. The source-word $\boldsymbol{\bar{Z}}$ is mapped to the code-word $\bar{\boldsymbol{z}}$ by $\bar{\boldsymbol{z}}=\boldsymbol{\bar{Z}}\boldsymbol{G}\>\left(\mathrm{mod}\>2\right)$. Note that any $K$ linearly independent code-words can be used to form the generating matrix. For an arbitrary $N_{v}$-dimensional binary vector $\boldsymbol{\bar{x}}=\left(\bar{x}_{1},\ldots,\bar{x}_{N_{v}}\right)\in\left\{0,1\right\}^{N_{v}}$, define the generalized syndrome vector $\boldsymbol{\bar{s}}\left(\boldsymbol{\bar{x}}\right)=\left(\bar{s}_{1}\left(\boldsymbol{\bar{x}}\right),\ldots,\bar{s}_{N_{c}}\left(\boldsymbol{\bar{x}}\right)\right)$ (generalized because it may not have $N_{v}-K$ bits) as $\boldsymbol{\bar{s}}\left(\boldsymbol{\bar{x}}\right)=\bar{\boldsymbol{x}}\boldsymbol{H}^{T}\>\left(\mathrm{mod}\>2\right)$. Then, $\bar{\boldsymbol{x}}$ is a code-word if and only if $\boldsymbol{\bar{s}}\left(\boldsymbol{\bar{x}}\right)=\boldsymbol{0}_{1\times N_{c}}$. This vector equation defines a set of generalized parity-check equations, consisting of a set of $N_{c}$ equations where only $N_{v}-K$ ones of them are linearly independent. It follows that we can choose the parity-check matrix for a given linear code in many ways and that we can define many syndrome vectors for the same code. The ratio of the length of the source-word to that of the code-word is called the rate: $r=K/N_{v}$. Later in this section, we will show specific examples of the matrices $\boldsymbol{G}$ and $\boldsymbol{H}$ for the PE architecture. The code-word $\bar{\boldsymbol{z}}$ is converted into a sequence of bipolar variables $\bar{\boldsymbol{z}}\rightarrow\boldsymbol{z}=\left(z_{1},\ldots,z_{N_{v}}\right)\in\left\{\pm 1\right\}^{N_{v}}$ (0 is mapped to +1+1, and $1$ to 1-1) and assumed to be modulated to antipodal signals $\left|v\right|\boldsymbol{z}$ by a binary phase shift keying modulation, where $\left|v\right|$ is the signal amplitude. We assume that the signal $\tilde{S}$ is transmitted over a channel subject to additive white Gaussian noise (AWGN). At the end of the channel, the receiver obtains an observation $\tilde{R}$, denoted by an antipodal vector $\boldsymbol{y}=\left(y_{1},\ldots,y_{N_{v}}\right)=\left|v\right|\boldsymbol{z}+\boldsymbol{n}\in\mathbb{R}^{N_{v}}$ , where $\boldsymbol{n}=\left(n_{1},\ldots,n_{N_{v}}\right)\in\mathbb{R}^{N_{v}}$ is a noise vector whose elements are i.i.d. Gaussian random variables with zero mean and variance $\sigma^{2}$. The purpose of ECC is to communicate reliably over a noisy channel. So, the goal of decoding is to reproduce the original source-word $\tilde{M}$ or associated code-word $\tilde{C}$ from the observation $\tilde{R}$.
@@
Figure 2: Two examples of a bipartite graph for $K=4$ logical spins. The dark blue circle with a label $\{k,l\}$ represents the spin variable $x_{kl}$, while the red circles with a label $\{k,l,m,n\}$ or $\{k,l,m\}$ represent the weight-4 syndrome $s_{klmn}^{(4)}$ and weight-3 syndrome $s_{klm}^{(3)}$, respectively. Let us relabel the variables with blue letters. An element of the code-word vector $\boldsymbol{x}=(x_{1},\ldots,x_{N_{v}})\in\{\pm 1\}^{N_{v}}$ is called a variable node (VN). The $i$-th syndrome of $\boldsymbol{x}$ is defined by $s_{i}(\boldsymbol{x})=\prod_{k\in N(i)}x_{k}\in\{\pm 1\}$ and an element of vector $\boldsymbol{s}(\boldsymbol{x})=(s_{1}(\boldsymbol{x}),\ldots,s_{N_{c}}(\boldsymbol{x}))\in\{\pm 1\}^{N_{c}}$ is called a check node (CN), where $N(i)=\{j:H_{ij}(H_{ij}^{{}^{\prime}})=1\}$ is the VNs adjacent to a CN $i$ $(1\leq i\leq N_{c})$ and $M(j)=\{i:H_{ij}(H_{ij}^{{}^{\prime}})=1\}$ is the CNs adjacent to a VN $j$ $(1\leq j\leq N_{v})$. The column and row weights of the parity-check matrix are defined by $d_{v}(i)=|M(j)|$ and $d_{c}(i)=|N(i)|$, respectively.
II.2 Parity encoding architecture and LDPC codes

### v4

@@ Abstract
We propose a practical hybrid decoding scheme for the parity-encoding architecture. This architecture was first introduced by N. Sourlas [Phys. Rev. Lett. 94, 070601 (2005)] as a computational technique for tackling hard optimization problems, especially those modeled by spin systems such as the Ising model and spin glasses, and reinvented by W. Lechner, P. Hauke, and P. Zoller [Sci. Adv., 1, e1500838 (2015)] to develop quantum annealing devices. We study the specific model, called the SLHZ model, aiming to achieve a near-term quantum annealing device implemented solely through geometrically local spin interactions. Taking account of the close connection between the SLHZ model and a classical low-density-parity-check code, two approaches can be chosen for the decoding: (1) finding the ground state of a spin Hamiltonian derived from the SLHZ model, which can be achieved via stochastic decoders such as a quantum annealer or a classical Monte Carlo sampler; (2) using deterministic decoding techniques for the classical LDPC code, such as belief propagation and bit-flip decoder. The proposed hybrid approach combines the two approaches by applying bit-flip decoding to the readout of the stochastic decoder based on the SLHZ model. We present simulations demonstrating that this approach can reveal the latent potential of the SLHZ model, realizing soft-annealing concept proposed by Sourlas.
@@ I INTRODUCTION
The spin-glass model is a fundamental concept in physics and mathematics, primarily used to study magnetism and phase transitions in statistical mechanics. A close connection between spin-glass models and combinatorial-optimization problems (COPs) is widely recognized and used as a tool for solving COPs [1]. The key idea is that the energy of the random Ising Hamiltonian can be viewed as a cost function the ground state of which corresponds to a solution of the COP. Probabilistic simulation of Ising spin dynamics, such as simulated annealing (SA) [2] and quantum annealing (QA) [3], is often used to find the ground state. The spin-glass model has applications in various industrial fields, including routing, scheduling, planning, decision-making, transportation, and telecommunications. A similar connection has been established between the spin-glass model and the classical error-correcting codes (ECCs). Sourlas formulated the decoding of the classical ECCs in terms of Bayesian inference [4]. He showed that the decoding can be expressed as a search for the ground state of the random Ising Hamiltonian. His study implies that decoding classical ECCs can be viewed as a COP, with the associated random Ising Hamiltonian corresponding to the cost function to be minimized (maximized).
@@
The complexity class of the COPs is known to be NP-hard. When we use the spin-glass models to tackle the COPs, the presence of many local minima separated by large energy barriers is a primary problem. The simulation algorithm gets trapped in those minima. To tackle this problem, Sourlas noted that two mathematically equivalent Hamiltonians can be derived for an ECC: a Hamiltonian given in terms of the original information source and an extended Hamiltonian given in terms of the information source and additional parity information [5, 6, 7]. The extended Hamiltonian involves extra parity-constraint terms. He proposed the soft-annealing concept, which uses the extended Hamiltonian as a new approach to solve hard COPs [8]. The key idea is that one can circumvent barriers and accelerate the dynamics of the algorithm by enlarging the space in which the problem is defined.
@@
On the other hand, the QA is expected to resolve the above problem through adiabatic quantum computing [9]. However, to apply a QA device to solve the COP universally, we need to simulate a fully connected graph model in a scalable manner using Ising spin hardware. The soft-annealing approach is also helpful for this purpose. Lechner, Hauke, and Zoller (LHZ) proposed utilizing this approach to embed a random Ising model with all-to-all connectivity within an enlarged Ising system with geometrically local interactions only [10]. They proposed the parity-encoding (PE) architecture and its concrete realization, which we refer to as the SLHZ model in this paper. In this model, Ising spins are arranged on a plane. They are connected by four-body interactions in a latticelike manner, which serve as a penalty that constrains the parities of spins. Both the local fields and couplings in the original spin-glass model are mapped to local fields that act on each spin in the SLHZ model. This architecture is particularly attractive for realizing near-term QA devices based on superconducting quantum bits [11, 12, 13, 14] because it is highly compatible with on-chip electronic circuitry [15, 16, 17, 18, 19]. LHZ also claimed that the SLHZ model is a programmable, scalable, and robust physical platform for QA. Notably, Sourlas was the first researcher to propose soft annealing using the SLHZ model and to recognize that this corresponds to the decoding of associated classical ECCs [8].
@@
Later, Pastawski and Preskill [20] studied the error-correcting capability of the SLHZ model. They noted that the soft annealing using the SLHZ model can be viewed as decoding of associated classical low-density parity-check (LDPC) code. Based on this observation, they applied the belief-propagation (BP) algorithm, a standard algorithm for LDPC codes [21], to decode the noisy readout of the SLHZ system [22, 23]. They found that one can obtain an error-free state from the noisy readout of the SLHZ model with a high probability if errors are generated by independent and identically distributed (IID) random spin-flip noise.
@@
To date, however, there has been little investigation focusing on whether the realistic classical decoding technique can fully exploit the potential of the SLHZ model. For example, Albash, Vincl, and Lidar have discussed whether several decoding algorithms may boost the performance of the QA device [24]. They have reported that simple majority-vote decoding (MVD) does not improve the performance of QA. Although they have found that more sophisticated minimum-weight decoding (MWD) can boost performance, it is NP-hard and an unrealistic approach for postreadout decoding of QA devices (for details, see Ref. [24] and Appendix A). In the context of developing near-term QA, it is preferable to devise a more practical decoding strategy tailored for the SLHZ model that can exploit its potential.
@@
This paper demonstrates that realistic classical decoding techniques for LDPC codes can unveil potential in the SLHZ model that has not been recognized to date. We propose a practical and straightforward postreadout decoding algorithm tailored to the SLHZ model. The proposed algorithm is an iterative hard-decision decoding algorithm based on majority voting of generalized syndromes, known as Gallager’s bit-flipping (BF) algorithm, which has been derived in the context of classical LDPC codes [22, 23]. Although the MWD is NP-hard, the BF algorithm is P, similar to the BP algorithms. Moreover, the BF algorithm requires less computational effort and processing time than the BP algorithm and can be implemented in digital logic. To demonstrate the BF algorithm, we have performed two simulations. First, assuming IID spin-flip noise, we have demonstrated that our BF decoding algorithm can recover the noisy output of the SLHZ model, eliminating spin-flip errors with performance comparable to that of the BP algorithm. Second, we have performed classical simulations of stochastically sampled spin readouts from the SLHZ model using a Markov-chain Monte Carlo (MCMC) sampler. This analysis allows us to investigate tolerance to leakage errors arising from dynamical and thermal excitations during the sampling. We show evidence that the BF decoding algorithm can efficiently eliminate spin-flip errors if we make the four-body penalty constraint sufficiently weak. This simulation suggests that a hybrid decoding approach, combining two algorithms—MCMC sampling followed by BF decoding—can be applied to a broader range of spin-flip noise and can mitigate the computational overhead of SLHZ systems. Our results are consistent with earlier work by Albash et al., who employed simulated quantum annealing to simulate the readout of the QA device based on the SLHZ model (hereafter referred to as the SLHZ-based QA device) when BP decoding was used. We try to understand this result by comparing our algorithm with several extended BF decoding algorithms. According to our analysis, this is a consequence of the combination of the Hamiltonian of the SLHZ system and the properties of BF or BP decoding. We believe that our BF decoding is promising for realizing a near-term SLHZ-based QA device.
@@
This paper is organized as follows. Section II explains the PE architecture, the SLHZ model, and the close connection between classical LDPC codes and the SLHZ model. Section III describes classical decoding of the SLHZ model based on Bayesian inference. We propose the BF algorithm as a simple, straightforward decoding algorithm tailored for the SLHZ model. Section IV demonstrates the proposed BF algorithm under the IID and non-IID spin-flip noise, and shows the performance. In Sec.V, we compare several BF decoding algorithms and discuss why our two-stage hybrid decoding strategy tolerates errors originating from non-IID spin-flip noise. We explain why two-stage hybrid decoding outperforms the two algorithms individually. Section VI concludes this paper. We also discuss the relevance and consistency of our study to the earlier work by Albash et al. [24] in Appendix A . Supplementary results are also provided in Appendix B .
II ERROR-CORRECTING CODES AND PARITY-ENCODING ARCHITECTURE
@@ II.1 Notion
In this paper, we assumed a discrete communication system (Fig.1). Suppose that a binary source word $\tilde{M}$ is first encoded into a binary code-word $\tilde{C}$ using some ECC. The code word is modulated into the physical signal $\tilde{S}$ and transmitted. The signal is affected by noise during transmission through a transmission channel. Let $\tilde{M}$ and $\tilde{C}$ be $K$ bits and $N_{v}\left(>K\right)$ bits, respectively, and denote them by the binary vectors $\bar{\textbf{Z}}=(\bar{Z}_{1},\ldots,\bar{Z}_{K})\in\{0,1\}^{K}$ and $\bar{\textbf{z}}=\left(\bar{z}_{1},\ldots,\bar{z}_{N_{v}}\right)\in\{0,1\}^{N_{v}}$, respectively. A linear code is defined by a one-to-one map from the set $\{\bar{\textbf{Z}}\}$ of $2^{K}$ source words $\tilde{M}$ of length $K$ to the set $\{\bar{\textbf{z}}\}$ of $2^{K}$ code words $\tilde{C}$ of length $N_{v}$. They are specified by a generating matrix $\textbf{G}_{K\times N_{v}}$ or a generalized parity check matrix $\textbf{H}_{N_{c}\times N_{v}}$ satisfying $\textbf{G}\textbf{H}^{T}=\boldsymbol{0}_{K\times N_{c}}\>\left(\mathrm{mod}\>2\right)$. Here G and H are binary matrices (i.e., their elements are 0 or $1$) and $\mathrm{mod}\>2$” denotes that the multiplication is modulo two. The source-word $\bar{\textbf{Z}}$ is mapped to the code word $\bar{\textbf{z}}$ by $\bar{\textbf{z}}=\bar{\textbf{Z}}\textbf{G}\>\left(\mathrm{mod}\>2\right)$. Note that any $K$ linearly independent code words can be used to form the generating matrix. For an arbitrary $N_{v}$-dimensional binary vector $\bar{\textbf{x}}=\left(\bar{x}_{1},\ldots,\bar{x}_{N_{v}}\right)\in\left\{0,1\right\}^{N_{v}}$, define the generalized syndrome vector $\bar{\textbf{s}}\left(\bar{\textbf{x}}\right)=\left(\bar{s}_{1}\left(\bar{\textbf{x}}\right),\ldots,\bar{s}_{N_{c}}\left(\bar{\textbf{x}}\right)\right)$ (generalized because it may not have $N_{v}-K$ bits) as $\bar{\textbf{s}}\left(\bar{\textbf{x}}\right)=\bar{\textbf{x}}\textbf{H}^{T}\>\left(\mathrm{mod}\>2\right)$. Then, $\bar{\textbf{x}}$ is a code word if and only if $\bar{\textbf{s}}\left(\bar{\textbf{x}}\right)=\boldsymbol{0}_{1\times N_{c}}$. This vector equation defines a set of generalized parity-check equations, consisting of a set of $N_{c}$ equations where only $N_{v}-K$ ones of them are linearly independent. It follows that we can choose the parity-check matrix for a given linear code in many ways and that we can define many syndrome vectors for the same code. The ratio of the length of the source word to that of the code word is called the rate: $r=K/N_{v}$. Later in this section, we will show specific examples of the matrices G and H for the PE architecture. The code word $\bar{\textbf{z}}$ is converted into a sequence of bipolar variables $\bar{\textbf{z}}\rightarrow\textbf{z}=\left(z_{1},\ldots,z_{N_{v}}\right)\in\left\{\pm 1\right\}^{N_{v}}$ (0 is mapped to +1+1, and $1$ to 1-1) and assumed to be modulated to antipodal signals $\left|v\right|\textbf{z}$ by a binary phase shift keying modulation, where $\left|v\right|$ is the signal amplitude. We assume that the signal $\tilde{S}$ is transmitted over a channel subject to additive white Gaussian noise (AWGN). At the end of the channel, the receiver obtains an observation $\tilde{R}$, denoted by an antipodal vector $\textbf{y}=\left(y_{1},\ldots,y_{N_{v}}\right)=\left|v\right|\textbf{z}+\textbf{n}\in\mathbb{R}^{N_{v}}$ , where $\textbf{n}=\left(n_{1},\ldots,n_{N_{v}}\right)\in\mathbb{R}^{N_{v}}$ is a noise vector :: the element of which are IID Gaussian random variables with zero mean and variance $\sigma^{2}$. The purpose of ECC is to communicate reliably over a noisy channel. So, the goal of decoding is to reproduce the original source word $\tilde{M}$ or associated code word $\tilde{C}$ from the observation $\tilde{R}$.
@@
Figure 2: Two examples of a bipartite graph for $K=4$ logical spins. The dark blue circle with a label $\{k,l\}$ represents the spin variable $x_{kl}$, while the red circles with a label $\{k,l,m,n\}$ or $\{k,l,m\}$ represent the weight-4 syndrome $s_{klmn}^{(4)}$ and weight-3 syndrome $s_{klm}^{(3)}$, respectively. Let us relabel the variables with blue letters. An element of the code-word vector $\textbf{x}=(x_{1},\ldots,x_{N_{v}})\in\{\pm 1\}^{N_{v}}$ is called a variable node (VN). The $i$ th syndrome of x is defined by $s_{i}(\textbf{x})=\prod_{k\in N(i)}x_{k}\in\{\pm 1\}$ and an element of vector $\textbf{s}(\textbf{x})=(s_{1}(\textbf{x}),\ldots,s_{N_{c}}(\textbf{x}))\in\{\pm 1\}^{N_{c}}$ is called a check node (CN), wheren the $N(i)=\{j:H_{ij}(H_{ij}^{{}^{\prime}})=1\}$ are the VNs adjacent to a CN $i$ $(1\leq i\leq N_{c})$ and the $M(j)=\{i:H_{ij}(H_{ij}^{{}^{\prime}})=1\}$ are the CNs adjacent to a VN $j$ $(1\leq j\leq N_{v})$. The column and row weights of the parity-check matrix are defined by $d_{v}(i)=|M(j)|$ and $d_{c}(i)=|N(i)|$, respectively. Table 1 shows the relevant parameters for your reference.
II.2 Parity-encoding architecture and LDPC codes